# Contextualized Policy Recovery: Modeling and Interpreting Medical Decisions with Adaptive Imitation Learning

### v3

@@ A.3 Training
    ADNI MRI scans MIMIC Antibiotics
    Class Algorithm $\lambda$ $k$ $\lambda$ $k$
    Interpretable CPR-RNN (ours) $0.0001$ $32$ $0.0001$ $32$
    CPR-LSTM (ours) $0.001$ $64$ $0.0001$ $32$
    Black-Box RNN - $64$ $32$
    LSTM - $64$ $64$
    Table 4: Hyperparameters chosen for different models.
@@ A.4.2 ADNI MRI scans
    ADNI MRI scans
    Class Algorithm AUROC AUPRC Brier $\downarrow$
    Interpretable Logistic regression $0.66\pm 0.03$ $0.86\pm 0.01$ $0.16\pm 0.01$
    INTERPOLE † $0.60\pm 0.04$ $0.81\pm 0.08$ $0.17\pm 0.05$
    INTERPOLE ‡ $0.44\pm 0.04$ $0.75\pm 0.09$ $0.19\pm 0.07$
    POETREE ‡ $0.62\pm 0.01$ $0.82\pm 0.01$ $0.18\pm 0.01$
    CPR-RNN (ours) $\mathbf{0.72\pm 0.02}$ $\mathbf{0.88\pm 0.02}$ $\mathbf{0.15\pm 0.01}$
    CPR-LSTM (ours) $\mathbf{0.72\pm 0.02}$ $\mathbf{0.88\pm 0.02}$ $\mathbf{0.15\pm 0.01}$
    Black-Box RNN $0.72\pm 0.02$ $0.88\pm 0.02$ $0.15\pm 0.01$
    LSTM $0.71\pm 0.02$ $0.88\pm 0.02$ $0.15\pm 0.01$
    Table 5: Action-matching performance of different policy learning algorithms on the ADNI MRI scans task. Bolded values in each column denote the best performance of any interpretable model. Listed performance of INTERPOLE and POETREE are from prior reports of state-of-the-art performance on these canonical datasets: † reported by Hüyük et al. (2022), ‡ reported by Pace et al. (2022).
@@
    Figure 8 shows how different contexts influence the agents decision function over time. Most notably, a high CDRSB value during the first two visits decreases the probability of ordering an MRI since this can be already seen as a strong indicator of dementia, making a scan less informative noa (2018). Patients A and B share the same decision model in $t=1$ since they both show medium CDRSB and avg hippocampal volume in $t=0$. Afterward, their decision function differs in $t=2$. Patient A’s hippocampal volume was measured as "low" in $t=1$ leading to a lower overall probability of ordering an MRI in $t=2$ indicated by a lower intercept and CDRSB coefficients. In contrast, Patient B was again diagnosed with an "avg" hippocampal volume in visit $t=1$. A low hippocampal volume can again be seen as a strong indicator of dementia making a scan less informative. Patient C, in contrast, is diagnosed with medium CDRSB and high hippocampal volume throughout all visits. This increases the probability of ordering an MRI (higher intercept) since there is no clear indication that would make an MRI obsolete.
@@ A.4.3 MIMIC Antibiotics
    MIMIC Antibiotics
    Class Algorithm AUROC AUPRC Brier $\downarrow$
    Interpretable Logistic regression $0.57\pm 0.03$ $0.80\pm 0.03$ $0.20\pm 0.01$
    INTERPOLE † NR NR NR
    INTERPOLE ‡ $0.65$ NR $0.21$
    POETREE ‡ $0.68$ NR $0.19$
    CPR-RNN (ours) $\mathbf{0.82\pm 0.01}$ $\mathbf{0.90\pm 0.01}$ $0.14\pm 0.01$
    CPR-LSTM (ours) $\mathbf{0.82\pm 0.01}$ $\mathbf{0.90\pm 0.01}$ $\mathbf{0.14\pm 0.00}$
    Black-Box RNN $0.83\pm 0.01$ $0.90\pm 0.01$ $0.13\pm 0.01$
    LSTM $0.84\pm 0.01$ $0.91\pm 0.01$ $0.13\pm 0.00$
    Table 6: Action-matching performance of different policy learning algorithms on the MIMIC antibiotics task. Bolded values in each column denote the best performance of any interpretable model. Listed performance of INTERPOLE and POETREE are from prior reports of state-of-the-art performance on these canonical datasets: † reported by Hüyük et al. (2022), ‡ reported by Pace et al. (2022). NR: No values reported.
@@
    Model Coefficients Figure 10 shows the heterogeneity in estimated model coefficients. The main drivers of heterogeneity are the intercept, potassium and creatinine coefficients.
    Treatment Trajectories To uncover typical treatment trajectories, we look at clusters of coefficients over 4 timesteps. We cluster patients into 5 subgroups using hierarchical clustering based on silhouette score ($0.566$) as seen in Figure 12. Figure LABEL:fig:coefficients_line shows the two largest clusters containing patients that get antibiotics throughout their stay (cluster $1$, $1707$ patients) and patients that never get antibiotics (cluster $5$, $384$ patients). The remaining three groups are plotted in Figure 11. Patients that get treatment for the first two days fall into cluster $3$ ($232$ patients). We can see that their treatment parametrization changes significantly after treatment is stopped. Cluster $4$ contains patients that get treated for the first day only ($240$ patients). Both patients in cluster $3$ and $4$ share similar decision models in $t=1$ since both were treated in $t=0$, it diverges in $t=2$ after treatment is stopped for one group and share the decision parametrization in $t=3$ after both groups were not treated in $t=2$. The remaining patients fall into cluster $4$ ($375$ patients).
    Appendix B Additional Plots

### v4

@@ A.4.2 ADNI MRI scans
ADNI MRI scans Class Algorithm AUROC AUPRC Brier $\downarrow$ Interpretable Logistic regression $0.66\pm 0.03$ $0.86\pm 0.01$ $0.16\pm 0.01$ INTERPOLE † $0.60\pm 0.04$ $0.81\pm 0.08$ $0.17\pm 0.05$ INTERPOLE ‡ $0.44\pm 0.04$ $0.75\pm 0.09$ $0.19\pm 0.07$ POETREE ‡ $0.62\pm 0.01$ $0.82\pm 0.01$ $0.18\pm 0.01$ CPR-RNN (ours) $\mathbf{0.72\pm 0.02}$ $\mathbf{0.88\pm 0.02}$ $\mathbf{0.15\pm 0.01}$ CPR-LSTM (ours) $\mathbf{0.72\pm 0.02}$ $\mathbf{0.88\pm 0.02}$ $\mathbf{0.15\pm 0.01}$ Black-Box RNN $0.72\pm 0.02$ $0.88\pm 0.02$ $0.15\pm 0.01$ LSTM $0.71\pm 0.02$ $0.88\pm 0.02$ $0.15\pm 0.01$
Table 5: Action-matching performance of different policy learning algorithms on the ADNI MRI scans task. Bolded values in each column denote the best performance of any interpretable model. Listed performance of INTERPOLE and POETREE are from prior reports of state-of-the-art performance on these datasets: † reported by (Hüyük et al., 2022), ‡ reported by (Pace et al., 2022).
@@
Figure 8 shows how different contexts influence the agents decision function over time. Most notably, a high CDRSB value during the first two visits decreases the probability of ordering an MRI since this can be already seen as a strong indicator of dementia, making a scan less informative (noa, 2018). Patients A and B share the same decision model in $t=1$ since they both show medium CDRSB and avg hippocampal volume in $t=0$. Afterward, their decision function differs in $t=2$. Patient A’s hippocampal volume was measured as "low" in $t=1$ leading to a lower overall probability of ordering an MRI in $t=2$ indicated by a lower intercept and CDRSB coefficients. In contrast, Patient B was diagnosed with an "avg" hippocampal volume in visit $t=1$. A low hippocampal volume can again be seen as a strong indicator of dementia making a scan less informative. Patient C, in contrast, is diagnosed with medium CDRSB and high hippocampal volume throughout all visits. This increases the probability of MRI (higher intercept) since there is no clear indication that would make an MRI obsolete.
@@ A.4.3 MIMIC Antibiotics
MIMIC Antibiotics Class Algorithm AUROC AUPRC Brier $\downarrow$ Interpretable Logistic regression $0.57\pm 0.03$ $0.80\pm 0.03$ $0.20\pm 0.01$ INTERPOLE † NR NR NR INTERPOLE ‡ $0.65$ NR $0.21$ POETREE ‡ $0.68$ NR $0.19$ CPR-RNN (ours) $\mathbf{0.82\pm 0.01}$ $\mathbf{0.90\pm 0.01}$ $0.14\pm 0.01$ CPR-LSTM (ours) $\mathbf{0.82\pm 0.01}$ $\mathbf{0.90\pm 0.01}$ $\mathbf{0.14\pm 0.00}$ Black-Box RNN $0.83\pm 0.01$ $0.90\pm 0.01$ $0.13\pm 0.01$ LSTM $0.84\pm 0.01$ $0.91\pm 0.01$ $0.13\pm 0.00$
Table 6: Action-matching performance of different policy learning algorithms on the MIMIC antibiotics task. Bolded values in each column denote the best performance of any interpretable model. Listed performance of INTERPOLE and POETREE are from prior reports of state-of-the-art performance on these canonical datasets: † reported by (Hüyük et al., 2022), ‡ reported by (Pace et al., 2022). NR: No values reported.
Model Coefficients Figure 12 shows the heterogeneity in estimated model coefficients. The main drivers of heterogeneity are the intercept, potassium and creatinine.
Treatment Trajectories To uncover typical treatment trajectories, we look at clusters of coefficients over 4 timesteps. We cluster patients into 5 subgroups using hierarchical clustering based on silhouette score ($0.566$) as seen in Figure 10. Figure 10 shows the two largest clusters containing patients that get antibiotics throughout their stay (cluster $1$, $1707$ patients) and patients that never get antibiotics (cluster $5$, $384$ patients). The remaining three groups are plotted in Figure 13. Patients that get treatment for the first two days fall into cluster $3$ ($232$ patients). We can see that their treatment parametrization changes significantly after treatment is stopped. Cluster $4$ contains patients that get treated for the first day only ($240$ patients). Both patients in cluster $3$ and $4$ share similar decision models in $t=1$ since both were treated in $t=0$, it diverges in $t=2$ after treatment is stopped for one group and share the decision parametrization in $t=3$ after both groups were not treated in $t=2$. The remaining patients fall into cluster $4$ ($375$ patients).
Global interpretability of CPR The coefficients $\beta$ of historical features and actions in previous time steps 0-3 and $\theta$ of current features at current time step 4 for predicting current treatment at time step 4 are shown in Fig. 11.
Figure 11: The magnitude of coefficients (linear effects) on each feature within CPR global policies at timestep 4 on MIMIC data. In CPR global, each of CPR’s predictions can be decomposed as a linear combination of all observed features and actions, rather than only interpreting current features.